Suppose the correlation coefficient is 0.80 and the number of observations is 62. What is the calculated test statistic? Is this significant correlation using a 1% level of significance? 2. Given that, r = 0.475 N = 10, Is this significant correlation between X and Y using a 5% level of significance? 


1
Expert's answer
2021-06-06T17:12:53-0400

"H0:\rho=0"

"Ha:\rho\ne0"

"t=\frac{r\sqrt{n-2}}{\sqrt{1-r^2}}"

"t=\frac{0.8\sqrt{60}}{\sqrt{1-0.8^2}}=10.33"

"Cv=t_{60,0.005}=-2.66, 2.66,"

10.33 falls in the rejection region. 10.33>2.66, this, we reject H0 and conclude that the correlation is significant.

2.

"H0:\rho=0"

"Ha:\rho\ne0"

"t=\frac{r\sqrt{n-2}}{\sqrt{1-r^2}}"

"t=\frac{0.475\sqrt{8}}{\sqrt{1-0.475^2}}=1.53"

"Cv=t_{8,0.025}=-2.306, 2.306."

1.5<2.306, we fail to reject the null hypothesis and conclude that there is no significant correlation between X and Y.